# Quasiconvexity in the Riemannian setting

### v2

@@ 2. Quasiconvexity in the Riemannian setting
For every $x\in M$ and $r>0$ smaller than the injectivity radius at $x$, we define the cube
@@
for every $\alpha,\beta\in\mathscr{L}(TU,\mathbb{R}^{m})$ such that $\alpha\in(T_{x}^{*}M)^{m}$ and $\beta\in(T_{y}^{*}M)^{m}$, where $d^{M}$ is the Riemannian distance on $M$. It is then easy to see that such distance, restricted to every fiber $(T_{x}^{*}M)^{m}$ coincides with the one associated to the “quadratic” norm (equivalent to the operatorial one) induced by the metric tensor $g$ of $M$, that we will denote with $|\cdot|_{g_{x}}$.
@@ 4. Some remarks and possible research directions
A natural continuation of our work would be to extend the previous results also to the $W^{1,p}(\Omega,\mathbb{R}^{m})$ setting, with $p\in[1,+\infty)$, in order to obtain the analogue of Theorem 1.3 (notice that the fact that the semicontinuity of the functional implies quasiconvexity of $f$ follows immediately by Theorem 3.2 in such setting, as in the Euclidean case).
Focardi and Spadaro in [4] showed an extension of the Acerbi–Fusco theorems (with $f$ continuous, like us) in the case of Sobolev maps $u:\Omega\to M$, with $\Omega$ an open bounded subset of $\mathbb{R}^{n}$ and $M$ a Riemannian manifold (this case actually “corresponds” to the Euclidean situation with a continuous function $f$ depending on $u$ and $Du$, but not on $x$). Therefore, a possible future research line could be to “combine” Theorem 3.3 with the results of Focardi and Spadaro, thus obtaining a characterization of quasiconvexity in the completely Riemannian context of Sobolev maps between two Riemannian manifolds.

### v3

@@ 2. Quasiconvexity in the Riemannian setting
For every $x\in M$ and $r>0$ such that $\sqrt{n\,}r/2$ is smaller than the injectivity radius of $M$ at $x$, we define the cube
@@
for every $\alpha,\beta\in\mathscr{L}(TU,\mathbb{R}^{m})$ such that $\alpha\in(T_{x}^{*}M)^{m}$ and $\beta\in(T_{y}^{*}M)^{m}$, where $d^{M}$ is the Riemannian distance on $M$. It is then easy to see that such distance, restricted to every fiber $(T_{x}^{*}M)^{m}$ coincides with the one associated to the “quadratic” norm (equivalent to the operator norm) induced by the metric tensor $g$ of $M$, that we will denote with $|\cdot|_{g_{x}}$.
@@ 4. Some remarks and possible research directions
A natural continuation of our work would be to extend the previous results to the $W^{1,p}(\Omega,\mathbb{R}^{m})$ setting, with $p\in[1,+\infty)$, in order to obtain the analogue of Theorem 1.3 (notice that the fact that the semicontinuity of the functional implies quasiconvexity of $f$ follows immediately by Theorem 3.2 in such setting, as in the Euclidean case).
Focardi and Spadaro in [4] showed an extension of the Acerbi–Fusco theorems (with $f$ continuous, like us) in the case of Sobolev maps $u:\Omega\to M$, with $\Omega$ an open bounded subset of $\mathbb{R}^{n}$ and $M$ a Riemannian manifold. Therefore, a possible future research line could be to “combine” Theorem 3.3 with the results of Focardi and Spadaro, thus obtaining a characterization of quasiconvexity in the completely Riemannian context of Sobolev maps between two Riemannian manifolds.